# An exact analytical solution for Dicke superradiance

### v4

@@ Abstract
We revisit the Dicke superradiance problem, namely finding the time evolution of the density operator for an initially inverted ensemble of $N$ identical two-level systems (spins), undergoing collective spontaneous decay. We derive a compactly written, easily computable and fully analytical solution, valid for any time during the evolution and for any $N$.
@@
$\dot{\rho}(t)=\mathcal{L}[\rho]=\Gamma\Bigg{[}S\rho S^{\dagger}-\frac{1}{2}\Big{(}S^{\dagger}S\rho-\rho S^{\dagger}S\Big{)}\Bigg{]}.$ (1)
@@
The solution proposed in Refs. [2, 3] has been mostly overlooked during the last decades; presumably the convoluted analytical expressions renders this solution computationally demanding, thus requiring resources similar to the direct simulation of the master equation, which is a trivial task for particles numbers of the order $N=10^{3}$ spins. The focus instead shifted more recently to deriving more informative solutions in the mesoscopic (or thermodynamic) limit of large $N$, such as, for example, in Refs. [6, 7, 8, 9, 10, 11]. A comprehensive early review of the problem and its implications can be found in Ref. [12].
While the original formulation of the Dicke superradiance model is impossible to realize, as it involves $N$ atoms spatially positioned at same point, a straightfoward realization can be achieved in cavity quantum electrodynamics, such as for example with atoms trapped in the standing wave field of a lossy optical cavity (with coupling and loss rates $g$ and $\kappa$, respectively). The cavity opens a decay channel with a collective collapse operator $S$ (after adiabatic elimination of the cavity mode) and a collapse rate $\Gamma=g^{2}/\kappa$. Assuming negligible independent spontaneous emission, the problem is described by Eq. (1). Such collective cavity superradiance effect has been recently implemented to advance superradiant lasers. Moreover, waveguide and circuit quantum electrodynamics provide alternative platforms to realize Dicke superradiance, where excited (artificial) atoms decay collectively into a nearby microwave or optical resonator mode.
@@
Figure 2: Distribution of Dicke states at the peak emission time, computed from the analytical expression in Eq. (3). The distrbution shows a more pronounced peak as the particle number is increased. The emission rate is computed as $I=\Gamma\langle S^{\dagger}S\rangle(t)$ and it peaks around $t_{\text{peak}}\approx\log{N}/(N\Gamma)$. We observe, that the distributions peak at $m=39$, $m=362$ and $m=3414$ respectively.
@@
Conclusions and outlook - We have provided a compact, fully analytical solution to the Dicke superradiance process, within the fully symmetric subspace of $N$ quantum emitters. The main observations have been i) the existence of simple recursive relations between Dicke states of index $m$ and $m-1$ (also previously observed in other works such as Refs. [11]), ii) the possibility to find a closed analytical solution for the recursive equation and iii) the identification of this solution as a sum of residues of a given simple function. This suggests that the solution could be expressed as a complex integral on a closed contour in the complex plane, which we will investigate in the future. We have also shown that the same simple formulation can be extended to the time evolution of any initial Dicke state as an initial state of the system.
The procedure followed here can be easily extended to provide fully analytical solutions to the case of the coherently driven Dicke superradiant system. Moreover, the analytical approach described here might provide ways to tackle the case of incoherently driven dissipative Dicke systems such is the case of superradiant lasing: in this case, $N$ quantum emitters are placed inside a lossy optical cavity or around optical fibers, where after the elimination of the cavity mode, the system behaves like a Dicke superradiant system with an external individual incoherent pump.

### v5

@@ Abstract
We revisit the Dicke superradiance problem, namely finding the time evolution of the density operator for an initially inverted ensemble of $N$ identical two-level systems (spins), undergoing collective spontaneous decay. We derive a compactly written, easily computable and fully analytical solution, valid at all times and $N$ during the evolution.
@@
$\dot{\rho}(t)=\mathcal{L}[\rho]=\Gamma\Bigg{[}S\rho S^{\dagger}-\frac{1}{2}\Big{(}S^{\dagger}S\rho+\rho S^{\dagger}S\Big{)}\Bigg{]}.$ (1)
@@
The solution proposed in Refs. [2, 3] has been mostly overlooked during the last decades; presumably the convoluted analytical expressions renders this solution computationally demanding, thus requiring resources similar to the direct simulation of the master equation, a trivial task only for particles numbers of the order $N=10^{3}$ spins. The focus instead shifted more recently to deriving more informative solutions in the mesoscopic (or thermodynamic) limit of large $N$, such as, for example, in Refs. [6, 7, 8, 9, 10, 11]. A comprehensive early review of the problem and its implications can be found in Ref. [12].
While the original formulation of the Dicke superradiance model in molecular gases is impossible to realize, as it involves $N$ two-level systems spatially positioned at same point, a straightfoward realization can be achieved in cavity quantum electrodynamics, such as for example with atoms trapped in the standing wave field of a lossy optical cavity (with coupling and loss rates $g$ and $\kappa$, respectively). The cavity opens a decay channel with a collective collapse operator $S$ (after adiabatic elimination of the cavity mode) and a collapse rate $\Gamma=g^{2}/\kappa$ [13]. Assuming negligible independent spontaneous emission, the problem is described by Eq. (1). Such collective cavity superradiance effect has been recently implemented to advance superradiant lasers. Moreover, waveguide and circuit quantum electrodynamics provide alternative platforms to realize Dicke superradiance, where excited (artificial) atoms decay collectively into a nearby microwave or optical resonator mode [14, 15].
@@
Figure 2: Distribution of Dicke states at the peak emission time, computed from the analytical expression in Eq. (3). The distrbution shows a more pronounced peak as the particle number is increased. The emission rate is computed as $I=\Gamma\langle S^{\dagger}S\rangle(t)$ and it peaks around $t_{\text{peak}}\approx\ln{N}/(N\Gamma)$. We observe, that the distributions peak at $m=39$, $m=362$ and $m=3414$ respectively.
@@
Conclusions and outlook - We have provided a compact, fully analytical solution to the Dicke superradiance process, within the fully symmetric subspace of $N$ two-level quantum emitters. The main observations have been i) the existence of simple recursive relations between Dicke states of index $m$ and $m-1$ (also previously observed in other works such as Refs. [11]), ii) the possibility to find a closed analytical solution for the recursive equation and iii) the identification of this solution as a sum of residues of a given simple function. This suggests that the solution could be expressed as a complex integral on a closed contour in the complex plane, provided in follow-up work [16]. We have also shown that the same simple formulation can be extended to the time evolution of any initial Dicke state as an initial state of the system.
The procedure followed here can be easily extended to provide fully analytical solutions to the case of the coherently driven Dicke superradiant system, which we plan to investigate in the future. Moreover, the analytical approach described here might provide ways to tackle the case of incoherently driven dissipative Dicke systems such is the case of superradiant lasing [13]: in this case, $N$ quantum emitters are placed inside a lossy optical cavity or around optical fibers, where after the elimination of the cavity mode, the system behaves like a Dicke superradiant system with an external individual incoherent pump.